### v1

# SinLlama - A Large Language Model for Sinhala

Abstract
Low-resource languages such as Sinhala are often overlooked by open-source Large Language Models (LLMs). In this research, we extend an existing multilingual LLM (Llama-3-8B) to better serve Sinhala. We enhance the LLM tokenizer with Sinhala specific vocabulary and perform continual pre-training on a cleaned 10 million Sinhala corpus, resulting in the SinLlama model. This is the very first decoder-based open-source LLM with explicit Sinhala support. When SinLlama was instruction fine-tuned for three text classification tasks, it outperformed base and instruct variants of Llama-3-8B by a significant margin.
Index Terms:
Sinhala, low-resource languages, large language models, continual pretraining, LLM, Llama, text classification
I Introduction
Large Language Models (LLMs) such as ChatGPT have significantly advanced Natural Language Processing (NLP) for languages with abundant training data and linguistic resources. While some recent successful open-source LLMs such as Llama [14] and Gemma [32] claim to have multilingual support, such multilingual LLMs (multiLLMs) continue to under-perform on low-resource languages (LRLs) [41]. Given that these LLMs are being created by large tech corporations, it is fair to assume that the decision on which languages to include in an LLM is driven by the perceived economic benefits in a global scale. As such, there is a risk of LRLs continue to be neglected, thus further increasing the digital divide between communities.
Sinhala is an Indo-Aryan language, which is only being used by a population of about 20 million in the island nation of Sri Lanka [10]. According to Ranathunga and De Silva [29]’s language categorization that uses the category definition of Joshi et al. [18], Sinhala is a low-resource language. There has been previous research that resulted in word embeddings and encoder-based language models for Sinhala [11]. However, techniques based on such models have become out-dated due to the demonstrated potential of LLMs. However, Sinhala, similar to many other LRLs, has not been considered during pre-training of any of the popular open-source LLMs such as Llama and Gemma released by the tech giants. While Sinhala is included in MaLA-500 [23] and Aya-101 [35] released by research teams, these models have not been able to outperform models such as Llama. On the other hand, MaLA-500 and Aya-101 models are large in size, incurring in large computer memory requirements.
To address these limitations, we present a focused adaptation of a state-of-the-art multiLLM to better support Sinhala. Our approach involves two key stages: (1) enhancing the model’s tokenizer by merging Sinhala-specific tokens into its vocabulary to improve subword segmentation; and (2) conducting continual pretraining using a 10 million unlabeled sentences corpus that includes cleaned and diverse Sinhala text, ensuring deeper linguistic representation.
We first carried out an investigation on the available open-source LLMs to select the best performing LLM for Sinhala. Llama-3 emerged as the most promising LLM. Then we further pre-trained the Llama-3-8B model as mentioned above. We term the resulting model, SinLlama. This model will be publicly released upon paper acceptance. In order to evaluate the performance of SinLlama against the original Llama-3, we fine-tuned both with data from three Sinhala benchmarks [37]: writing-style classification, news categorization and sentiment analysis. The results show that when fine-tuned on Sinhala task-specific data, SinLlama significantly outperforms both fine-tuned Llama-base and Llama-instruct versions.
II Related Work
II-A Large Language Models (LLMs)
Pre-trained Language Models (PLMs), created using the Transformer architecture [36], have significantly advanced the NLP landscape. Among them, decoder-only models such as ChatGPT have become especially popular due to their strong performance. Open-source large-scale decoder models, known as LLMs (e.g., LLaMA [14], Mistral [17], Gemma [32]), are trained on much larger datasets and offer performance comparable to commercial systems. Training an LLM has two primary stages: training a tokenizer and self-supervised training with an unlabeled raw text corpus. The latter step is termed pre-training, and the resulting model is termed a base model. However, such base models lack task-specific capabilities. Therefore, these base models should be fine-tuned with task-specific instruction data, which is also termed instruct-tuning. The aforementioned open-source LLMs are available in both base and insturct versions.
II-B Empirical Studies on LLM Performance for LRLs
The evaluation of LLMs across multiple languages reveals significant performance disparities between high-resource and LRLs. Ahuja et al. [2]’s MEGA framework conducted the first comprehensive benchmarking of LLMs on 16 NLP tasks across 70 diverse languages (including Sinhala) for comparing the performance of different models on XLSUM dataset. They showed that while models such as GPT-4 performed well in high-resource languages, they struggle with LRLs, especially those with unique morphological structures or scripts.
Recent surveys of LLMs underscore the influence of pretraining data, language typology, and evaluation benchmarks on model performance across languages. Chang et al. [4] conducted a comprehensive review of over 200 LLMs, highlighting that multilingual performance is significantly affected by the diversity and representativeness of the pretraining corpus, with underrepresented languages, especially those with complex morphology or distinct scripts, often exhibiting degraded results. The survey also emphasizes the limitations of current multilingual benchmarks, which are either narrow in task scope or limited in linguistic coverage, leading to an incomplete understanding of LLM capabilities in low-resource languages.
II-C Extending LLMs for New Languages
There are several strategies to extend an LLM to an unseen language.
  • Continual Pre-training: This involves further training an existing LLM on text data from the target language to improve its language understanding and generation capabilities. Continued pre-training can improve alignment with the new language, but may risk degrading performance in previously learned languages if not carefully managed [7].
  • Vocabulary Extension: Adding new tokens specific to the target language to the tokenizer’s vocabulary improves encoding efficiency and model performance. Simple vocabulary extension combined with continued pre-training has been shown to be effective in various languages [33].
  • Fine-tuning: Instruction tuning or task-specific fine-tuning on data in the new language helps the model adapt to language-specific nuances and tasks. Techniques such as translation-assisted chain-of-thought fine-tuning (TaCo) [34] have been proposed to improve cross-lingual transfer for LRLs.
  • Preference Optimization: This method optimizes the model based on human preference data, which can be scarce for LRLs. Preference optimization helps align model output more closely with human expectations in the new language context.
  • In-context Learning: Providing relevant examples to LLMs during inference is known as in-context learning or few-shot learning. In-context learning has shown to work in the context of LRLs as well [25].
  • Leveraging Lexicons: Structured language-specific linguistic resources such as lexicons can be integrated with LLMs to boost their performance for LRLs [19]. Unlike zero-shot methods, which often fail in domain-specific settings, few-shot prompts enhance both accuracy and efficiency, marking a shift toward more adaptive strategies.
II-D Sinhala Language Computing
While there has been research on implementing Sinhala language computing tools, most of has been done using rule-based, statistical or early-generation deep learning techniques [10]. In the NLP field, these techniques have long become obsolete. Nevertheless, there have been some attempts to keep Sinhala NLP up-to-date. Lakmal et al. [21] presented Word2Vec and FastText embeddings for Sinhala and carried out a detailed investigation. Dhananjaya et al. [11] built the encoder-based language model SinBERT, by fine-tuning an encoder-based Transformer [24] with 15.7 million Sinhala sentences.
III Model Selection
In the initial phase of our research, we aimed to identify the most suitable open-source multiLLM, focusing on LRLs. The first round of filtration was done referring to empirical research that compared open-source LLM. This resulted in the following models: Llama-3-8B [14], MaLA-500 [23], Aya-101 [35], PolyLM [38], XGLM, [22] Falcon [3], BLOOMZ [26], Phi3 [1], Mistral [17], OPT [40], Aya-23 [35], and BigTrans [39]. These models are listed in Table I. Based on criteria including multilingual capability, parameter size (fewer than 8 billion parameters) [^1] Model size has a direct relationship with the memory requirement in using an LLM., Sinhala language representation, and performance metrics, we selected Llama-3-8B [14] for further experimentation.
TABLE I: Model Details and Sinhala Language Support
Model Parameters Official Sinhala Support No:of languages
Llama-3-8B 8B No 8
MaLA-500 10.7B Yes 534
Aya-101 13B Yes 101
PolyLM-13B 13B No 18
XGLM-7.5B 7.5B Yes 30
Falcon-7B 7B No 15
BLOOMZ-7.1B 7.1B No 46
Phi-3-mini 3.8B No 14
Mistral-7B 7B No 10
OPT-66B 66B No 5
Aya-23-8B 8B No 23
BigTrans-13B 13B Yes 102
Llama-3-8B [14] officially supports 8 languages and is pre-trained on a large dataset over 15 trillion data tokens, sourced from publicly available Internet materials. It has 128K context length. Lalama-3-8B outperformed its counterparts such as Mistral 7B and Gemma 7B on common benchmarks such as MMLU [15], ARC [6], DROP [12] and GPQA [31].
IV Dataset
IV-A Pre-training Data
Pre-training of an LLM has to be done with a substantially large dataset, in order to guarantee it learns the language-specific information. Therefore, we created a dataset by combining MADLAD-400 [20] and CulturaX [27] for the Sinhala language [^2] The dataset created by Dhananjaya et al. [11] is no longer accessible. MADLAD-400 is a manually audited, general-domain 3 Trillion token monolingual dataset based on CommonCrawl [28], spanning 419 languages. CulturaX is a multilingual dataset with 6.3 Trillion tokens in 167 languages. Although both have been cleaned, there were some entries that contained non-Sinhala words. Because of that, heuristics-based filtering [13] was applied to remove such noisy sentences. These include rules such as minimum sentence length, non-Sinhala sentence removal, exclusion of sentences with irregular punctuation, and elimination of URLs and numeric prefixes.For the pre-training, we combined both datasets and removed duplicates. Our final dataset contains 10,730,154 (10.7M) Sinhala sentences, comprising 303,958,959 (303.96M) tokens. This dataset will be released upon paper acceptance.
IV-B Fine-tuning Data
For the supervised fine-tuning experiments, we focused on Sinhala text classification tasks. We selected three publicly available datasets: News Category Classification [9], Sentiment Analysis [30], and Writing Style Classification [11].These datasets collectively represent a diverse range of linguistic features and styles, making them suitable for evaluating the effectiveness of LLMs for Sinhala.
Since the original datasets were not in a directly usable format for fine-tuning, several preprocessing steps were performed. The Sentiment classification dataset, originally in XML format, required extraction of relevant fields: content within the <phrase> tag was used as the input text, and the <sentiment> tag provided the corresponding label. For other datasets, we only applied standard text cleaning procedures such as deduplication, removal of malformed entries, and whitespace normalization.
To enable robust evaluation, we manually split each dataset into training, validation, and test sets using an 80:10:10 ratio. We employed stratified sampling to ensure that each split preserved the original label distribution, thereby preventing skewed performance metrics. The final dataset sizes used for fine-tuning are shown in Table II.
TABLE II: Dataset sizes for Text Classificiation Task
Dataset Train Test Validation
Sentiment Tagger[30] 10010 1252 1252
Writting style classification[11] 7238 905 905
Sinhala news category classification[9] 2661 333 333
V Continual Pretraining to Build SinLlama
We used vocabulary extension and continuous pre-training (see Section II-C) on Llama-3-8B to create SinLlama. We trained a tokenizer using the pre-training corpus using tiktoken [^3] https://github.com/openai/tiktoken, and merged this tokenizer with the Llama 3 tokenizer. We used this extended tokenizer and performed continual pretraining using the codebase from Chinese-Llama [8]. The only modification made to the original Chinese-Llama hyperparameters was the reduction of the block_size from 1024 to 512444Larger block sizes cause memory allocation issues on the GPUs we used, preventing successful model loading..
VI Task-Specific Finetuning
VI-A Prompt Selection
For both inference and supervised fine-tuning, it was essential to select the most effective prompt templates from the available datasets. We adopted the prompt format introduced by the Alpaca Instruct Template [5] as in Fig.  1. To identify the most suitable prompts for each classification task, we referred to the work by Ahuja et al. [2], which provided well-designed prompts for several tasks.
Figure 1: Prompt template for supervised fine-tuning
VI-B Fine-tuning
TABLE III: Performance Comparison Across Tasks
Model Writing Style News Category Sentiment Tagger
Precision Recall F1 Precision Recall F1 Precision Recall F1
Llama-3-8B base 33.091 20.927 24.504 23.904 18.919 19.031 41.942 38.011 36.285
Llama-3-8B base - Finetuned 71.060 38.179 49.451 64.770 63.664 61.136 61.110 61.215 59.353
Llama-3-8B instruct 50.782 50.719 48.755 31.866 30.330 29.219 48.451 36.796 31.257
Llama-3-8B instruct - Finetuned 70.996 35.942 42.256 50.883 50.450 47.812 68.856 68.729 68.784
SinLlama 33.315 4.393 7.531 44.166 19.520 15.614 47.153 25.856 22.689
SinLlama - Finetuned 85.906 52.157 58.893 89.033 86.787 86.402 75.246 70.055 72.471
We fine-tuned SinLlama and the Llama-3-8B base model separately with the training split of each dataset mentioned in Section IV-B, creating a set of task-specific fine-tuned models. We used LoRA fine-tuning [16], which enables faster training and lower memory usage.
Figure 2 shows an example experimental setup for one dataset; we followed a similar process for other datasets.
Figure 2: An example experimental setup for the Writing Style Classification dataset.
VII Evaluation
We used Llama-3-8B base original model, Llama-3-8B-instruct original model, and their fine-tuned (with our data) versions, as the baselines. From the results of table III, we can identify that the fine-tuned SinLlama model performs best at each of the task with a significantly high margin. When the Llama-3-8B base model is fine-tuned with our dataset, there is a significant gain over the base model. However, this result is far below the fine-tuned SinLlama model. Fine-tuned Llama-3-8B instruct model also outperforms its un-fine-tuned version. However, it falls below the fine-tuned Llama-3-8B base model, which is surprising. Overall, the writing style classification task seems to be the most difficult for the LLMs. Because the input field of that dataset is much larger than those of the other two datasets.
VIII Conclusion
Although Sinhala has been included in several multilingual pre-trained large language models, it was not intentionally targeted during the pre-training phase. In this study, we introduce SinLlama, the first-ever large language model pre-trained specifically for Sinhala. Our results demonstrate that fine-tuning SinLlama with Sinhala task-specific data results in significant gains compared to fine-tuning Llama-3-8B base and instruct models. In the future, we plan to conduct experiments on several fine-tuning strategies and preference optimization methods.
Appendix A Prompts Used
All of the prompts used for the benchmark process referred to the prompts given from the Ahuja et al. [2]. Selected prompts are given below.
A-A Writing-style-classification
Feature columns in the dataset are ”comments”, ”labels”.
You are an NLP assistant whose purpose is to classify
Sinhala comments into predefined categories. The
categories are as follows: ACADEMIC, CREATIVE, NEWS,
BLOG. Given a Sinhala comment, your task is to
determine which category it belongs to. Choose one
category: ACADEMIC, CREATIVE, NEWS, or BLOG. Answer
must match exactly in capitalization and formatting.
Comment: {comment}
Answer:
A-B Sentiment-tagger
The original dataset consist of XML tags. After the processing we created a dataset with the following features. Feature columns in the dataset are ”context”, ”text”, ”sentiment”.
You are an NLP assistant whose purpose is to solve
Sentiment Analysis problems. Sentiment Analysis is
the task of determining whether the sentiment,
opinion or emotion expressed in a textual data is.
Does the following Sinhala sentence have a POSITIVE,
NEGATIVE or NEUTRAL sentiment? {text}. Give only
the POSITIVE, NEGATIVE, or NEUTRAL as the answer in
one word.
A-C Sinhala-News-Category-classification
Feature columns in the dataset are ”comments”, ”labels”.
You are an NLP assistant whose purpose is to classify
Sinhala comments into predefined categories. The
categories are as follows: (Political: 0, Business: 1,
Technology: 2, Sports: 3, Entertainment: 4) Given a
Sinhala comment, your task is to determine which
category it belongs to. Please only give the label
as a number (0, 1, 2, 3, or 4) that best represents
the comments category. comment: {text} answer:

### v2

@@
 Low-resource languages such as Sinhala are often overlooked by open-source Large Language Models (LLMs).
-In this research, we extend an existing multilingual LLM (Llama-3-8B) to better serve Sinhala.
-We enhance the LLM tokenizer with Sinhala specific vocabulary and perform continual pre-training on a cleaned 10 million Sinhala corpus, resulting in the SinLlama model.
+Therefore, it is imperative that the existing LLMs are further trained to cover such languages.
+In this research, we extend an existing multilingual LLM (Llama-3-8B) to get a better coverage for Sinhala.
+We enhanced the LLM tokenizer with Sinhala specific vocabulary and performed continual pre-training on a 10 million sentence Sinhala corpus, resulting in the SinLlama model.
 This is the very first decoder-based open-source LLM with explicit Sinhala support.
@@
 I Introduction
-Large Language Models (LLMs) such as ChatGPT have significantly advanced Natural Language Processing (NLP) for languages with abundant training data and linguistic resources.
-While some recent successful open-source LLMs such as Llama [14] and Gemma [32] claim to have multilingual support, such multilingual LLMs (multiLLMs) continue to under-perform on low-resource languages (LRLs) [41].
+Large Language Models (LLMs) such as ChatGPT have significantly advanced the field of Natural Language Processing (NLP) for languages with abundant training data and linguistic resources.
+However, their support for low-resource languages (LRLs) is significantly low.
+While some recent successful open-source LLMs such as Llama [1] and Gemma [2] claim to have multilingual support, such multilingual LLMs (multiLLMs) continue to under-perform on LRLs [3].
 Given that these LLMs are being created by large tech corporations, it is fair to assume that the decision on which languages to include in an LLM is driven by the perceived economic benefits in a global scale.
@@
 While Sinhala is included in MaLA-500 [8] and Aya-101 [9] released by research teams, these models have not been able to outperform models such as Llama.
-On the other hand, MaLA-500 and Aya-101 models are large in size, incurring in large computer memory requirements.
+On the other hand, MaLA-500 and Aya-101 models are large in size, incurring in large computer memory requirements when used for downstream tasks.
 To address these limitations, we present a focused adaptation of a state-of-the-art multiLLM to better support Sinhala.
-Our approach involves two key stages: (1) enhancing the model’s tokenizer by merging Sinhala-specific tokens into its vocabulary to improve subword segmentation; and (2) conducting continual pretraining using a 10 million unlabeled sentences corpus that includes cleaned and diverse Sinhala text, ensuring deeper linguistic representation.
+Our approach involves two key stages: (1) enhancing the model’s tokenizer by merging Sinhala-specific tokens into its vocabulary to improve subword segmentation; and (2) conducting continual pretraining using a 10 million sentence corpus that includes cleaned and diverse Sinhala text, ensuring deeper linguistic representation.
 We first carried out an investigation on the available open-source LLMs to select the best performing LLM for Sinhala.
@@
 Then we further pre-trained the Llama-3-8B model as mentioned above.
-We term the resulting model, SinLlama.
-This model will be publicly released upon paper acceptance.
-In order to evaluate the performance of SinLlama against the original Llama-3, we fine-tuned both with data from three Sinhala benchmarks [37]: writing-style classification, news categorization and sentiment analysis.
+We term the resulting model, SinLlama [^1] https://huggingface.co/polyglots/SinLlama_v01, which is publicly released.
+In order to evaluate the performance of SinLlama against the original Llama-3, we fine-tuned both with data from three Sinhala benchmark datasetss [10]: writing-style classification, news categorization and sentiment analysis.
 The results show that when fine-tuned on Sinhala task-specific data, SinLlama significantly outperforms both fine-tuned Llama-base and Llama-instruct versions.
@@
 Pre-trained Language Models (PLMs), created using the Transformer architecture [11], have significantly advanced the NLP landscape.
-Among them, decoder-only models such as ChatGPT have become especially popular due to their strong performance.
-Open-source large-scale decoder models, known as LLMs (e.g., LLaMA [14], Mistral [17], Gemma [32]), are trained on much larger datasets and offer performance comparable to commercial systems.
+Among them, decoder-only models such as ChatGPT have become especially popular due to their strong performance across a multitude of tasks.
+Open-source large-scale decoder models, known as LLMs (e.g., LLaMA, Mistral [12], Gemma), are also trained on larger datasets and offer performance comparable to commercial systems.
 Training an LLM has two primary stages: training a tokenizer and self-supervised training with an unlabeled raw text corpus.
@@
 However, such base models lack task-specific capabilities.
-Therefore, these base models should be fine-tuned with task-specific instruction data, which is also termed instruct-tuning.
-The aforementioned open-source LLMs are available in both base and insturct versions.
+Therefore, these base models should be fine-tuned with task-specific instruction data, which is termed fine-tuning or instruct-tuning.
+The aforementioned open-source LLMs are available in both base and instruct versions.
 II-B Empirical Studies on LLM Performance for LRLs
@@
 [13]’s MEGA framework conducted the first comprehensive benchmarking of LLMs on 16 NLP tasks across 70 diverse languages (including Sinhala) for comparing the performance of different models on XLSUM dataset.
-They showed that while models such as GPT-4 performed well in high-resource languages, they struggle with LRLs, especially those with unique morphological structures or scripts.
+They showed that while models such as GPT-4 performed well in high-resource languages, they struggle with LRLs, especially those with unique morphological structures and/or scripts.
 Recent surveys of LLMs underscore the influence of pretraining data, language typology, and evaluation benchmarks on model performance across languages.
@@
 II-D Sinhala Language Computing
-While there has been research on implementing Sinhala language computing tools, most of has been done using rule-based, statistical or early-generation deep learning techniques [10].
+While there has been research on implementing Sinhala language computing tools, most of it has been done using rule-based, statistical or early-generation deep learning techniques [4].
 In the NLP field, these techniques have long become obsolete.
@@
 In the initial phase of our research, we aimed to identify the most suitable open-source multiLLM, focusing on LRLs.
-The first round of filtration was done referring to empirical research that compared open-source LLM.
+The first round of was carried out referring to empirical research that compared open-source LLMs.
 This resulted in the following models: Llama-3-8B [1], MaLA-500 [8], Aya-101 [9], PolyLM [22], XGLM, [23] Falcon [24], BLOOMZ [25], Phi3 [26], Mistral [12], OPT [27], Aya-23 [9], and BigTrans [28].
@@
 TABLE I: Model Details and Sinhala Language Support
-Model Parameters Official Sinhala Support No:of languages
+Model Parameters
+Official
+Sinhala Support
+Number of
+languages
 Llama-3-8B 8B No 8
@@
 [7] is no longer accessible.
+According to Ranathunga et al.
+[35], this is a common problem in LRL domain..
 MADLAD-400 is a manually audited, general-domain 3 Trillion token monolingual dataset based on CommonCrawl [36], spanning 419 languages.
@@
 Our final dataset contains 10,730,154 (10.7M) Sinhala sentences, comprising 303,958,959 (303.96M) tokens.
-This dataset will be released upon paper acceptance.
+This dataset is publicly released [^4] https://huggingface.co/datasets/polyglots/MADLAD_CulturaX_cleaned.
 IV-B Fine-tuning Data
@@
 Dataset Train Test Validation
-Sentiment Tagger[30] 10010 1252 1252
+Sentiment analysis [39] 10010 1252 1252
 Writting style classification [7] 7238 905 905
@@
 We used this extended tokenizer and performed continual pretraining using the codebase from Chinese-Llama [40].
-The only modification made to the original Chinese-Llama hyperparameters was the reduction of the block_size from 1024 to 512444Larger block sizes cause memory allocation issues on the GPUs we used, preventing successful model loading..
+The only modification made to the original Chinese-Llama hyperparameters was the reduction of the block_size from 1024 to 512666Larger block sizes cause memory allocation issues on the GPUs we used, preventing successful model loading..
 VI Task-Specific Finetuning
@@
 In the future, we plan to conduct experiments on several fine-tuning strategies and preference optimization methods.
+Acknowledgment
+We thank CSIR - Central Scientific Instruments Organization, India, and Emojot (Pvt) Ltd for providing the computing resources.
 Appendix A Prompts Used
@@
 A-A Writing-style-classification
-Feature columns in the dataset are ”comments”, ”labels”.
+Feature columns in the dataset are “comments”, “labels”.
 You are an NLP assistant whose purpose is to classify
@@
 A-B Sentiment-tagger
-The original dataset consist of XML tags. After the processing we created a dataset with the following features. Feature columns in the dataset are ”context”, ”text”, ”sentiment”.
+The original dataset consist of XML tags. After the processing we created a dataset with the following features. Feature columns in the dataset are “context”, “text”, “sentiment”.
 You are an NLP assistant whose purpose is to solve
@@
 A-C Sinhala-News-Category-classification
-Feature columns in the dataset are ”comments”, ”labels”.
+Feature columns in the dataset are “comments”, “labels”.
 You are an NLP assistant whose purpose is to classify